OpenStudy (anonymous):

Rectangle A has an area of 9 - x2. Rectangle B has an area of x2 + x - 12. In simplest form, what is the ratio of the area of Rectangle A to the area of Rectangle B? Show your work.

OpenStudy (anonymous):

divide the area of triangle A/B \[\frac{ 9-x ^{2} }{ x ^{2}+x-12 }\]

OpenStudy (anonymous):

now simplify

OpenStudy (anonymous):

\[\frac{ (3-x)(3+x) }{ (x+4)(x-3) }\]

OpenStudy (anonymous):

okay umm

OpenStudy (anonymous):

\[\frac{ (3-x)(3+x) }{ -(x+4)(3-x)}\]

OpenStudy (anonymous):

@Abdulhameed You just flip it around?

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

well i looked at it and notice u switched the x and 3 on the bottom right and made a negative off the side on the bottom left

OpenStudy (anonymous):

\[-\frac{ (3+x) }{(x+4) }\]

OpenStudy (anonymous):

so thats it?

OpenStudy (anonymous):

\[\frac{ -x-3 }{ x+4 }\]

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

oh okay ty

OpenStudy (anonymous):

do u think u can help me with one more?

OpenStudy (anonymous):

u r welcome , ok i'll do my best to help

OpenStudy (anonymous):

Trevor is tiling his bathroom floor, which has an area that is represented as 117r4 square inches. Each tile has an area of square root of the quantity 9 r to the thirteenth power . The total number of tiles used can be represented by the expression below. one hundred seventeen r to the fourth power, all over the square root of the quantity nine r to the thirteenth power

OpenStudy (anonymous):

and what do i have to find in the question ??

OpenStudy (anonymous):

Simplify the expression for the total number of tiles used.

OpenStudy (anonymous):

@Abdulhameed

OpenStudy (anonymous):

\[\frac{ 117r ^{4} }{ \sqrt{9r ^{13}} }\]

OpenStudy (anonymous):

\[\sqrt{9}=3\] so \[\frac{ 117r ^{4} }{ 3\times \sqrt{r ^{13}}} \]

OpenStudy (anonymous):

117/3=39 \[\frac{ 39r ^{4} }{ r ^{13/2} }\] \[39r ^{4-13/2}\] \[39r ^{-5/2}\] \[\frac{ 39 }{ r ^{5/2} }\] \[\frac{ 39 }{ \sqrt{r ^{5}}}\]

OpenStudy (anonymous):

dat all? o:

OpenStudy (anonymous):

yep that's all :)

OpenStudy (anonymous):

Ty so much

OpenStudy (anonymous):

u can skip some steps bcoz i wrote some steps to show what i did (for explanation)

OpenStudy (anonymous):

alright ty

OpenStudy (anonymous):

np :)